12(2k+11)=12(2k+12)
Answer by stanbon(75887)

12(2k+11)=12(2k+12)
----
Divide both sides by 12 to get:
2k + 11 = 2k + 12
-----
Subtract 2k from both sides:
11 = 12
----
That contradiction implies the original equation
is a contradiction. There are no solutions.
====================
Cheers,
Stan H.
==============